1.     Suppose market demand is given Qd=a-bp as and market supply is given as, Qs=c+dp and then find the following

o  Equilibrium price

o  Equilibrium quantity

o  Price elasticity of demand at equilibrium

o  Price elasticity of supply at equilibrium 


1
Expert's answer
2021-11-23T11:05:45-0500

Solution:

a.). At equilibrium: Qd = Qs

a – bp = c + dp

p = "c - \frac{a}{b} + d"

Equilibrium price = "c - \frac{a}{b} + d"

 

b.). Equilibrium quantity:

Plug price into the demand function to derive the equilibrium quantity:

Qd = a – bp

Qd = a – b("c - \frac{a}{b} + d" )

Qd = "2ab + ad - \frac{cb}{b} + d"

Equilibrium quantity: "2ab + ad - \frac{cb}{b} + d"


c.). Price elasticity of demand at equilibrium:

Price elasticity of demand (PED) = "\frac{\triangle Qd}{P} \times \frac{P}{Qd}"

"\frac{\triangle Qd}{P} = -b"


PED = "-b \times \frac{c-a}{b+d} \div \frac{2ab+ad-cb}{b+d}"


PED = "-b\frac{c-a}{2ab+ad-cb}"

 

d.). Price elasticity of supply at equilibrium:

Price elasticity of supply (PES) = "\frac{\triangle Qs}{P} \times \frac{P}{Qs}"

"\frac{\triangle Qs}{P} = d"

PES = "d \times \frac{c-a}{b+d} \div \frac{cb+2cd-da}{b+d}"

 

PES ="\frac{d(c-a)}{cb+2cd - da}"